A stunt driver gives her car 3600 N to take a jump that has an angle of inclination of 31°. If she and the car together weigh 4700 N, what is the acceleration of the car? (Assume no friction.)


1
Expert's answer
2021-02-25T11:25:16-0500

"mg=4700N"

"\Rightarrow m=\dfrac{4700}{9.8}=479.59" kg


Balancing the forces along the plane,

"ma=F-4700sin31^{\omicron}"

"\Rightarrow a=\dfrac{3600}{479.59}-\dfrac{4700}{479.59}sin31^{\omicron}"

"\Rightarrow a=2.554m/s^2"